# Sliced Rényi Pufferfish Privacy: Directional Additive Noise Mechanism and Private Learning with Gradient Clipping

### v2

@@ Appendix D Minimality and Attainability of HUC
Step 2. By Definition 5.3, the given sequence $h^{\mathsf{ms}}=\{h^{\mathsf{ms}}_{t}\}_{t=1}^{T}$ with $h^{\mathsf{ms}}_{t}=(h^{\mathsf{ms}}_{t,1},\dots,h^{\mathsf{ms}}_{t,m})$ is a valid ms-HUC vector if, for every $\theta\in\Theta$, $(s_{i},s_{j})\in\mathcal{Q}$, history $y_{<t}$ in the support, and coupling $\gamma\in\Pi(\mu_{s_{i}}^{\theta},\mu_{s_{j}}^{\theta})$, we have, for $\ell=1,\dots,m$,
$\mathbb{E}_{\circledR}\mathbb{E}_{((X,X^{\prime}),R_{t})\sim\gamma\times\mathbb{P}_{\eta,\rho}}\bigl[\bigl|\Delta_{t,\ell}(X,X^{\prime};\circledR)\bigr|^{2}\bigr]\leq h^{\mathsf{ms}}_{t,\ell}.$
In the SRPP–SGD construction of Proposition 5.2, we have
$h_{t,\ell}(\circledR)=\Bigl(\frac{2L_{t,\ell}C}{B_{t}}\Bigr)^{2}K_{t}(\circledR)^{2},$
for a random discrepancy $K_{t}(\circledR)$, and
$h^{\mathsf{ms}}_{t,\ell}=\Bigl(\frac{2L_{t,\ell}C}{B_{t}}\Bigr)^{2}\overline{K}_{t}^{2},$
where $\overline{K}_{t}^{2}$ is any scalar satisfying $\overline{K}_{t}^{2}\geq\mathbb{E}[K_{t}(\circledR)^{2}]$.
Therefore
$\mathbb{E}_{\circledR}\bigl[h_{t,\ell}(\circledR)\bigr]=\Bigl(\frac{2L_{t,\ell}C}{B_{t}}\Bigr)^{2}\mathbb{E}_{\circledR}\bigl[K_{t}(\circledR)^{2}\bigr]\leq\Bigl(\frac{2L_{t,\ell}C}{B_{t}}\Bigr)^{2}\overline{K}_{t}^{2}=h^{\mathsf{ms}}_{t,\ell}$,
and similarly
$\mathbb{E}_{\circledR}\bigl[\max_{\ell}h_{t,\ell}(\circledR)\bigr]\leq\max_{\ell}h^{\mathsf{ms}}_{t,\ell}.$
[…]
@@
$\mathbb{E}_{\circledR}\Bigl[\mathtt{AveSD}^{\omega}_{\alpha}\bigl(\mathcal{M}^{\theta}_{\circledR}(s_{i})\,\|\,\mathcal{M}^{\theta}_{\circledR}(s_{j})\bigr)\Bigr]\leq\frac{\alpha}{2\sigma^{2}}\sum_{t=1}^{T}\sum_{\ell=1}^{m}\omega_{\ell}\mathbb{E}_{\circledR}[h_{t,\ell}(\circledR)].$
Using $\mathbb{E}_{\circledR}[h_{t,\ell}(\circledR)]\leq h^{\mathsf{ms}}_{t,\ell}$, we obtain
@@
$\mathbb{E}_{\circledR}\Bigl[\mathtt{AveSD}^{\omega}_{\alpha}\bigl(\mathcal{M}^{\theta}_{\circledR}(s_{i})\,\|\,\mathcal{M}^{\theta}_{\circledR}(s_{j})\bigr)\Bigr]\leq\frac{\alpha}{2\sigma^{2}}sum_{t=1}^{T}\sum_{\ell=1}^{m}\omega_{\ell}h^{\mathsf{ms}}_{t,\ell}.$
@@
$\sigma^{2}\geq\frac{\alpha}{2\epsilon}\sum_{t=1}^{T}\sum_{\ell=1}^{m}\omega_{\ell}h^{\mathsf{ms}}_{t,\ell}$,
then
$\mathbb{E}_{\circledR}\Bigl[\mathtt{AveSD}^{\omega}_{\alpha}\bigl(\mathcal{M}^{\theta}_{\circledR}(s_{i})\,\|\,\mathcal{M}^{\theta}_{\circledR}(s_{j})\bigr)\Bigr]\leq\epsilon$
for all $\theta\in\Theta$ and $(s_{i},s_{j})\in\mathcal{Q}$.
By Definition 5.4, this is exactly the $(\alpha,\epsilon,\omega)$-ms-Ave-SRPP guarantee in part (i).
Step 4. Similarly, taking expectations in (47) gives
$\mathbb{E}_{\circledR}\Bigl[\mathtt{JSD}^{\omega}_{\alpha}\!\bigl(\mathcal{M}^{\theta}_{\circledR}(s_{i})\,\|\,\mathcal{M}^{\theta}_{\circledR}(s_{j})\bigr)\Bigr]\leq\frac{\alpha}{2\sigma^{2}}\sum_{t=1}^{T}\mathbb{E}_{\circledR}\bigl[\max_{\ell}h_{t,\ell}(\circledR)\bigr].$
[…]
Thus
$\max_{\ell}h_{t,\ell}(\circledR)=\Bigl(\max_{\ell}a_{t,\ell}\Bigr)K_{t}(\circledR)^{2},$
and
$\mathbb{E}_{\circledR}\bigl[\max_{\ell}h_{t,\ell}(\circledR)\bigr]=\Bigl(\max_{\ell}a_{t,\ell}\Bigr)\mathbb{E}_{\circledR}\bigl[K_{t}(\circledR)^{2}\bigr]$
$\leq\max_{\ell}a_{t,\ell}\,\overline{K}_{t}^{2}=\max_{\ell}h^{\mathsf{ms}}_{t,\ell}$,
where we used $\overline{K}_{t}^{2}\geq\mathbb{E}[K_{t}(\circledR)^{2}]$ and the closed form $h^{\mathsf{ms}}_{t,\ell}=a_{t,\ell}\overline{K}_{t}^{2}$ from Proposition 5.2.
Hence
$\mathbb{E}_{\circledR}\Bigl[\mathtt{JSD}^{\omega}_{\alpha}\bigl(\mathcal{M}^{\theta}_{\circledR}(s_{i})\,\|\,\mathcal{M}^{\theta}_{\circledR}(s_{j})\bigr)\Bigr]\leq\frac{\alpha}{2\sigma^{2}}\sum_{t=1}^{T}\max_{\ell}h^{\mathsf{ms}}_{t,\ell}.$
@@
$\sigma^{2}\geq\frac{\alpha}{2\epsilon}\sum_{t=1}^{T}\max_{\ell}h^{\mathsf{ms}}_{t,\ell}$,
then
$\mathbb{E}_{\circledR}\Bigl[\mathtt{JSD}^{\omega}_{\alpha}\bigl(\mathcal{M}^{\theta}_{\circledR}(s_{i})\,\|\,\mathcal{M}^{\theta}_{\circledR}(s_{j})\bigr)\Bigr]\leq\epsilon$
for all $\theta\in\Theta$ and $(s_{i},s_{j})\in\mathcal{Q}$, which is exactly the $(\alpha,\epsilon,\omega)$-ms-Joint-SRPP guarantee in part (ii) by Definition 5.5. ∎
Appendix N Proof of Theorem 5.8
@@
Case 1: $\Xi\in\{\text{Ave},\text{ms-Ave}\}$. Fix a slice set $\mathcal{U}=\{u_{\ell}\}_{\ell=1}^{m}\subset\mathbb{S}^{d-1}$. For each mechanism $\mathcal{M}_{j}$, slice $u_{\ell}$, and dataset realization $x$, define the 1-D projected probability measure
$\mathrm{P}_{j,\ell}^{x}=\Pr\bigl(\langle\mathcal{M}_{j}(x),u_{\ell}\rangle\in\cdot\,\big|\,s_{i},\theta\bigr),$
$\mathrm{Q}_{j,\ell}^{x}=\Pr\bigl(\langle\mathcal{M}_{j}(x),u_{\ell}\rangle\in\cdot\,\big|\,s_{j},\theta\bigr).$
Likewise, for the composed mechanism and fixed $x$,
$\mathrm{P}^{\mathrm{comp},x}_{\ell}=\Pr\bigl(\langle\vec{\mathcal{M}}(x),u_{\ell}\rangle\in\cdot\,\big|\,s_{i},\theta\bigr),$
$\mathrm{Q}^{\mathrm{comp},x}_{\ell}=\Pr\bigl(\langle\vec{\mathcal{M}}(x),u_{\ell}\rangle\in\cdot\,\big|\,s_{j},\theta\bigr).$
From the independence of the primitive randomness across $j$ (for this fixed $x$), we have
$\mathrm{P}^{\mathrm{comp},x}_{\ell}=\bigotimes_{j=1}^{J}\mathrm{P}_{j,\ell}^{x}\quad\textup{and}\quad\mathrm{Q}^{\mathrm{comp},x}_{\ell}=\bigotimes_{j=1}^{J}\mathrm{Q}_{j,\ell}^{x}.$
Applying (48) to these 1-D projections gives, for every $x$ and $\ell$,
$\mathtt{D}_{\alpha}\bigl(\mathrm{P}^{\mathrm{comp},x}_{\ell}\big\|\mathrm{Q}^{\mathrm{comp},x}_{\ell}\bigr)=\sum_{j=1}^{J}\mathtt{D}_{\alpha}(\mathrm{P}_{j,\ell}^{x}\|\mathrm{Q}_{j,\ell}^{x}).$ (49)
By the definition of $(\alpha,\epsilon_{j},\omega)$-Ave-SRPP, if $\mathcal{M}_{j}$ is $(\alpha,\epsilon_{j},\omega)$-Ave-SRPP then, for every prior $\theta$, secret pair $(s_{i},s_{j})$, and data belief (coupling), the $\omega$-weighted average of the sliced Rényi costs of $\mathcal{M}_{j}$ is at most $\epsilon_{j}$. In particular, this bound holds uniformly for every dataset realization $x$ in the support of $\mu_{s_{i}}^{\theta}$ and $\mu_{s_{j}}^{\theta}$, so we have
$\sum_{\ell=1}^{m}\omega_{\ell}\mathtt{D}_{\alpha}(\mathrm{P}_{j,\ell}^{x}\|\mathrm{Q}_{j,\ell}^{x})\leq\epsilon_{j},$
for all $x,(s_{i},s_{j}),\theta$.
For the composed mechanism and fixed $d$, the average sliced Rényi divergence is
$\mathtt{AveSD}_{\omega}^{\alpha}\bigl(\mathrm{P}^{\mathrm{comp},x}\|\mathrm{Q}^{\mathrm{comp},x}\bigr)=\sum_{\ell=1}^{m}\omega_{\ell}\mathtt{D}_{\alpha}\bigl(\mathrm{P}^{\mathrm{comp},x}_{\ell}\big\|\mathrm{Q}^{\mathrm{comp},x}_{\ell}\bigr)$
@@
$\leq\sum_{j=1}^{J}\epsilon_{j}.$
Since this bound holds for every dataset realization $x$ and for all $\theta\in\Theta$ and $(s_{i},s_{j})\in\mathcal{Q}$, taking the supremum over all admissible beliefs and couplings shows that the composed mechanism $\vec{\mathcal{M}}$ is $(\alpha,\sum_{j=1}^{J}\epsilon_{j},\omega)$-Ave-SRPP.
For ms-Ave-SRPP, the same argument applies verbatim after folding the subsampling randomness into the mechanism’s output: ms-Ave-SRPP is just Ave-SRPP applied to the channel that includes both SGD updates and subsampling randomness. Thus $\vec{\mathcal{M}}$ is also $(\alpha,\sum_{j=1}^{J}\epsilon_{j},\omega)$-ms-Ave-SRPP whenever each $\mathcal{M}_{j}$ is $(\alpha,\epsilon_{j},\omega)$-ms-Ave-SRPP.
Case 2: $\Xi\in\{\text{Joint},\text{ms-Joint}\}$. By Definition 4.3 and the discussion in Appendix B, Joint-SRPP can be interpreted as standard Rényi Pufferfish Privacy applied to the sliced channel
$\mathtt{Slice}\circ\mathcal{M}:X\longmapsto(U,Z)\quad\text{with}\quad U\sim\omega,Z=\langle\mathcal{M}(X),U\rangle.$
More precisely, for any mechanism $\mathcal{M}$ we have
@@
$\mathtt{JSD}^{\omega}_{\alpha}\bigl(\Pr(\mathcal{M}(X)\mid s_{i},\theta)\,\big\|\,\Pr(\mathcal{M}(X)\mid s_{j},\theta)\bigr)$
$=\mathtt{D}_{\alpha}\Bigl(\Pr(\mathtt{Slice}\circ\mathcal{M}(X)\mid s_{i},\theta)\,\Big\|\,\Pr(\mathtt{Slice}\circ\mathcal{M}(X)\mid s_{j},\theta)\Bigr).$
@@
Thus, if $\mathcal{M}_{j}$ is $(\alpha,\epsilon_{j},\omega)$-Joint-SRPP, then the sliced channel $\mathtt{Slice}\circ\mathcal{M}_{j}$ is $(\alpha,\epsilon_{j})$-RPP in the underlying Pufferfish scenario.
Consider the sliced channels
$\mathcal{N}_{j}:=\mathtt{Slice}\circ\mathcal{M}_{j},\quad j=1,\dots,J,$
and their composition
$\vec{\mathcal{N}}(X)=\bigl(\mathcal{N}_{1}(X),\dots,\mathcal{N}_{J}(X)\bigr)$
@@
$=\bigl(\mathtt{Slice}\circ\mathcal{M}_{1}(X),\dots,\mathtt{Slice}\circ\mathcal{M}_{J}(X)\bigr).$
By independence of the primitive randomness across $j$, the same process of (48) implies that standard 1-D RPP enjoys additive composition: if each $\mathcal{N}_{j}$ is $(\alpha,\epsilon_{j})$-RPP, then $\vec{\mathcal{N}}$ is $(\alpha,\sum_{j=1}^{J}\epsilon_{j})$-RPP. Equivalently, for every $\theta$ and $(s_{i},s_{j})$,
@@
$\mathtt{D}_{\alpha}\Bigl(\Pr\bigl(\vec{\mathcal{N}}(X)\mid s_{i},\theta\bigr)\,\Big\|\,\Pr\bigl(\vec{\mathcal{N}}(X)\mid s_{j},\theta\bigr)\Bigr)\leq\sum_{j=1}^{J}\varepsilon_{j}.$
Finally, observe that applying the sliced channel $\mathtt{Slice}$ to the composed mechanism $\vec{\mathcal{M}}$ produces exactly $\vec{\mathcal{N}}$ as a post-processing:
$\vec{\mathcal{N}}=\mathtt{Slice}\circ\vec{\mathcal{M}}.$
Since Rényi divergence (and therefore 1-D RPP) is non-increasing under post-processing, the above bound on $\vec{\mathcal{N}}$ implies the same bound for the Joint-SRD of $\vec{\mathcal{M}}$, i.e.
$\mathtt{JSD}^{\omega}_{\alpha}\bigl(\Pr(\vec{\mathcal{M}}(x)\mid s_{i},\theta)\,\big\|\,\Pr(\vec{\mathcal{M}}(x)\mid s_{j},\theta)\bigr)\leq\sum_{j=1}^{J}\epsilon_{j}.$
Thus, by taking the supremum over all $\theta$ and $(s_{i},s_{j})$, we can show that $\vec{\mathcal{M}}$ is $(\alpha,\sum_{j=1}^{J}\epsilon_{j},\omega)$-Joint-SRPP. As in the average case, ms-Joint-SRPP is defined by applying Joint-SRPP to the channel that includes subsampling randomness in its output. Hence the same composition argument applies directly, giving $(\alpha,\sum_{j=1}^{J}\epsilon_{j},\omega)$-ms-Joint-SRPP whenever each $\mathcal{M}_{j}$ is $(\alpha,\epsilon_{j},\omega)$-ms-Joint-SRPP.
@@
Appendix O Experiment Details

### v3

@@ Appendix H Minimality and Attainability of HUC
Step 1. Fix an iteration $t\in\{1,\dots,T\}$ and a slice $u_{\ell}\in\mathcal{U}$. Fix any history $y_{<t}$ in the support and any coupling $\gamma\in\Pi(\mu^{\theta}_{s_{i}},\mu^{\theta}_{s_{j}})$. Apply Lemma U.1 with $u=u_{\ell}$: for $\gamma$-a.e. $(X,X^{\prime})$,
$\mathbb{E}_{R_{t}}\!\left[\mathtt{D}_{\alpha}\!\left(\Pr(\langle Y_{t},u_{\ell}\rangle\mid s_{i},\theta,R_{t})\ \Big\|\ \Pr(\langle Y_{t},u_{\ell}\rangle\mid s_{j},\theta,R_{t})\right)\right]$
$=\frac{\alpha}{2\sigma^{2}}\,\mathbb{E}_{R_{t}}\!\left[\big|\langle f_{t}(X,y_{<t};R_{t})-f_{t}(X^{\prime},y_{<t};R_{t}),u_{\ell}\rangle\big|^{2}\right].$
Since $h_{t}^{\mathsf{sa}}$ is a valid sa-HUC vector (Definition 5.3), the expectation on the right-hand side is at most $h^{\mathsf{sa}}_{t,\ell}$. Therefore,
$\mathbb{E}_{R_{t}}\!\left[\mathtt{D}_{\alpha}\!\left(\Pr(\langle Y_{t},u_{\ell}\rangle\mid s_{i},\theta,R_{t})\ \Big\|\ \Pr(\langle Y_{t},u_{\ell}\rangle\mid s_{j},\theta,R_{t})\right)\right]$ (72)
$\leq\frac{\alpha}{2\sigma^{2}}\,h^{\mathsf{sa}}_{t,\ell}.$
Step 2 (sa-Ave-SRPP). Fix a realization $\mathbf{r}=(r_{1},\dots,r_{T})$ of the subsampling outcomes $\mathbf{R}$. Apply the same conditional SRPP accountant (composition argument) as used for Theorem 5.6 to the conditional mechanism $\mathcal{M}^{\theta,\mathbf{r}}$. This yields the pointwise (in $\mathbf{r}$) bound
@@
$\mathtt{AveSD}^{\omega}_{\alpha}\!\left(\mathcal{M}^{\theta,\mathbf{r}}_{s_{i}}\ \Big\|\ \mathcal{M}^{\theta,\mathbf{r}}_{s_{j}}\right)\leq\sum_{t=1}^{T}\sum_{\ell=1}^{m}\omega_{\ell}\,$ (73)
$\mathtt{D}_{\alpha}\!\left(\Pr(\langle Y_{t},u_{\ell}\rangle\mid s_{i},\theta,R_{t}=r_{t})\ \Big\|\ \Pr(\langle Y_{t},u_{\ell}\rangle\mid s_{j},\theta,R_{t}=r_{t})\right).$
Take expectation over $\eta$ and use linearity and the fact that the $t$-th summand depends only on $\eta_{t}$:
@@
$\mathbb{E}_{\eta}\!\left[\mathtt{AveSD}^{\omega}_{\alpha}\!\left(\mathcal{M}^{\theta,\eta}_{s_{i}}\ \Big\|\ \mathcal{M}^{\theta,\eta}_{s_{j}}\right)\right]\leq\sum_{t=1}^{T}\sum_{\ell=1}^{m}\omega_{\ell}\,\mathbb{E}_{\eta_{t}}\!\Big[$
$\mathtt{D}_{\alpha}\!\left(\Pr(\langle Y_{t},u_{\ell}\rangle\mid s_{i},\theta,R_{t})\ \Big\|\ \Pr(\langle Y_{t},u_{\ell}\rangle\mid s_{j},\theta,R_{t})\right)\Big].$
Applying (72) yields
$\mathbb{E}_{\eta}\!\left[\mathtt{AveSD}^{\omega}_{\alpha}\!\left(\mathcal{M}^{\theta,\eta}_{s_{i}}\ \Big\|\ \mathcal{M}^{\theta,\eta}_{s_{j}}\right)\right]\leq\frac{\alpha}{2\sigma^{2}}\sum_{t=1}^{T}\sum_{\ell=1}^{m}\omega_{\ell}\,h^{\mathsf{sa}}_{t,\ell}.$
@@
$\sigma^{2}\geq\frac{\alpha}{2\epsilon}\sum_{t=1}^{T}\sum_{\ell=1}^{m}\omega_{\ell}\,h^{\mathsf{sa}}_{t,\ell}$,
then the right-hand side is at most $\epsilon$. Since $\theta$ and $(s_{i},s_{j})$ were arbitrary, Definition 5.4 implies that Algorithm 1 satisfies $(\alpha,\epsilon,\omega)$-sa-Ave-SRPP, proving (i).
Step 3 (sa-Joint-SRPP). Fix a realization $\mathbf{r}=(r_{1},\dots,r_{T})$ of the subsampling outcomes $\mathbf{R}$. Apply the same conditional accountant as in Step 2 for the joint sliced divergence:
$\mathtt{JSD}^{\omega}_{\alpha}\!\left(\mathcal{M}^{\theta,\mathbf{r}}_{s_{i}}\ \Big\|\ \mathcal{M}^{\theta,\mathbf{r}}_{s_{j}}\right)\leq\sum_{t=1}^{T}\max_{\ell\in[m]}\,\mathtt{D}_{\alpha}\!\Big(\Pr(\langle Y_{t},u_{\ell}\rangle\mid s_{i},\theta,R_{t}=r_{t})\$(74)
$\Big\|\ \Pr(\langle Y_{t},u_{\ell}\rangle\mid s_{j},\theta,R_{t}=r_{t})\Big),$
Take expectation over $r_{t}$. and use linearity across $t$:
$\mathbb{E}_{\eta}\!\left[\mathtt{JSD}^{\omega}_{\alpha}\!\left(\mathcal{M}^{\theta,\eta}_{s_{i}}\ \Big\|\ \mathcal{M}^{\theta,\eta}_{s_{j}}\right)\right]\leq\sum_{t=1}^{T}\mathbb{E}_{\eta_{t}}\!\Big[$
$\max_{\ell\in[m]}\,\mathtt{D}_{\alpha}\!\left(\Pr(\langle Y_{t},u_{\ell}\rangle\mid s_{i},\theta,R_{t})\ \Big\|\ \Pr(\langle Y_{t},u_{\ell}\rangle\mid s_{j},\theta,R_{t})\right)\Big].$
Now invoke the explicit SRPP–SGD sa-HUC instantiation from Proposition 5.2. In that setting, for each $t$ and $\ell$ there is a deterministic coefficient
$a_{t,\ell}:=\Bigl(\frac{2L_{t,\ell}C}{B_{t}}\Bigr)^{2}$
and a (scalar) discrepancy random variable $K_{t}(R_{t})$ (shared across slices at round $t$) such that for $\gamma$-a.e. $(X,X^{\prime})$ and all $\ell$,
$\big|\langle f_{t}(X,y_{<t};R_{t})-f_{t}(X^{\prime},y_{<t};R_{t}),u_{\ell}\rangle\big|^{2}\leq a_{t,\ell}\,K_{t}(R_{t})^{2}.$
Condition on $R_{t}=r$ and apply the same Gaussian identity as in Lemma U.1 to get
$\mathtt{D}_{\alpha}\!\left(\Pr(\langle Y_{t},u_{\ell}\rangle\mid s_{i},\theta,R_{t}=r)\ \Big\|\ \Pr(\langle Y_{t},u_{\ell}\rangle\mid s_{j},\theta,R_{t}=r)\right)$
$=\frac{\alpha}{2\sigma^{2}}\,\big|\langle f_{t}(X,y_{<t};r)-f_{t}(X^{\prime},y_{<t};r),u_{\ell}\rangle\big|^{2}$
$\leq\frac{\alpha}{2\sigma^{2}}\,a_{t,\ell}\,K_{t}(r)^{2}.$
Taking the maximum over $\ell$ yields, for each $r$,
$\max_{\ell\in[m]}\mathtt{D}_{\alpha}(\cdots\mid R_{t}=r)\leq\frac{\alpha}{2\sigma^{2}}\,\Bigl(\max_{\ell\in[m]}a_{t,\ell}\Bigr)\,K_{t}(r)^{2}.$
Now take expectation over $R_{t}$:
$\mathbb{E}_{R_{t}}\!\left[\max_{\ell\in[m]}\mathtt{D}_{\alpha}(\cdots)\right]\leq\frac{\alpha}{2\sigma^{2}}\,\Bigl(\max_{\ell\in[m]}a_{t,\ell}\Bigr)\,\mathbb{E}_{R_{t}}\!\left[K_{t}(R_{t})^{2}\right].$
By the mean-square discrepancy cap (17), $\mathbb{E}_{R_{t}}[K_{t}(R_{t})^{2}]\leq\overline{K}_{t}^{2}$, hence
$\mathbb{E}_{R_{t}}\!\left[\max_{\ell\in[m]}\mathtt{D}_{\alpha}(\cdots)\right]\leq\frac{\alpha}{2\sigma^{2}}\,\Bigl(\max_{\ell\in[m]}a_{t,\ell}\Bigr)\,\overline{K}_{t}^{2}=\frac{\alpha}{2\sigma^{2}}\,\max_{\ell\in[m]}h^{\mathsf{sa}}_{t,\ell}$,
where the last equality uses the closed form $h^{\mathsf{sa}}_{t,\ell}=a_{t,\ell}\overline{K}_{t}^{2}$ from Proposition 5.2. Substituting back,
$\mathbb{E}_{\eta}\!\left[\mathtt{JSD}^{\omega}_{\alpha}\!\left(\mathcal{M}^{\theta,\eta}_{s_{i}}\ \Big\|\ \mathcal{M}^{\theta,\eta}_{s_{j}}\right)\right]\leq\frac{\alpha}{2\sigma^{2}}\sum_{t=1}^{T}\max_{\ell\in[m]}h^{\mathsf{sa}}_{t,\ell}.$
@@
$\sigma^{2}\geq\frac{\alpha}{2\epsilon}\sum_{t=1}^{T}\max_{\ell\in[m]}h^{\mathsf{sa}}_{t,\ell}$,
then the right-hand side is at most $\epsilon$. Since $\theta$ and $(s_{i},s_{j})$ were arbitrary, Definition 5.5 implies that Algorithm 1 satisfies $(\alpha,\epsilon,\omega)$-sa-Joint-SRPP, proving (ii). ∎
Appendix V Proof of Theorem 5.8
@@
Case 1: $\Xi\in\{\textup{Ave},\textup{ms-Ave}\}$. Fix $\theta\in\Theta$ and $(s_{i},s_{j})\in\mathcal{Q}$ with $P_{\theta}^{S}(s_{i}),P_{\theta}^{S}(s_{j})>0$. Fix any coupling $\gamma\in\Pi(\mu_{s_{i}}^{\theta},\mu_{s_{j}}^{\theta})$ and take $(X,X^{\prime})\sim\gamma$. For each $j$, write $W_{j}:=\mathcal{M}_{j}(X)$ and $W^{\prime}_{j}:=\mathcal{M}_{j}(X^{\prime})$.
Fix a slice $u_{\ell}\in\mathcal{U}$ and consider the 1-D release sequence
$Z^{(\ell)}_{1:J}=\big(\langle W_{1},u_{\ell}\rangle,\dots,\langle W_{J},u_{\ell}\rangle\big),$
${Z^{\prime}}^{(\ell)}_{1:J}=\big(\langle W^{\prime}_{1},u_{\ell}\rangle,\dots,\langle W^{\prime}_{J},u_{\ell}\rangle\big).$
Let $\mathrm{P}^{(\ell)}:=\Pr(Z^{(\ell)}_{1:J}\in\cdot\mid s_{i},\theta)$ and $\mathrm{Q}^{(\ell)}:=\Pr({Z^{\prime}}^{(\ell)}_{1:J}\in\cdot\mid s_{j},\theta)$.
Per-step conditional slice bounds. Because each $\mathcal{M}_{j}$ is an SRPP-SGD mechanism whose SRPP guarantee is obtained from an HUC/sa-HUC sequence (Theorem 5.6 or Theorem 5.7), its per-slice Rényi bound is uniform over all histories and (coupled) dataset pairs. In particular, for every $j\in[J]$, every $\ell\in[m]$, and every history $z^{(\ell)}_{<j}$ in the support, we have
@@
$\big\|\ \Pr(\langle W^{\prime}_{j},u_{\ell}\rangle\in\cdot\mid s_{j},\theta,{Z^{\prime}}^{(\ell)}_{<j}=z^{(\ell)}_{<j})\Big)\leq\varepsilon_{j,\ell},$
for some nonnegative numbers $\{\varepsilon_{j,\ell}\}_{\ell=1}^{m}$ satisfying
$\sum_{\ell=1}^{m}\omega_{\ell}\,\varepsilon_{j,\ell}\leq\epsilon_{j}.$ (77)
(Here $\varepsilon_{j,\ell}$ is the per-slice budget induced by the HUC/sa-HUC-based bound for $\mathcal{M}_{j}$.)
[…]
$\mathtt{D}_{\alpha}(\mathrm{P}^{(\ell)}\|\mathrm{Q}^{(\ell)})\leq\sum_{j=1}^{J}\varepsilon_{j,\ell}.$
Now average over slices:
@@
$\mathtt{AveSD}^{\omega}_{\alpha}\bigl(\Pr(\vec{\mathcal{M}}(X)\mid s_{i},\theta)\,\big\|\,\Pr(\vec{\mathcal{M}}(X)\mid s_{j},\theta)\bigr)$
$=\sum_{\ell=1}^{m}\omega_{\ell}\,\mathtt{D}_{\alpha}(\mathrm{P}^{(\ell)}\|\mathrm{Q}^{(\ell)})\leq\sum_{\ell=1}^{m}\omega_{\ell}\sum_{j=1}^{J}\varepsilon_{j,\ell}$
$=\sum_{j=1}^{J}\sum_{\ell=1}^{m}\omega_{\ell}\varepsilon_{j,\ell}\leq\sum_{j=1}^{J}\epsilon_{j},$
where the last inequality uses (77). Since $\theta$ and $(s_{i},s_{j})$ were arbitrary, this proves $(\alpha,\sum_{j=1}^{J}\epsilon_{j},\omega)$-Ave-SRPP for $\vec{\mathcal{M}}$.
For ms-Ave-SRPP, the same argument applies verbatim after folding the subsampling randomness into each mechanism’s output (i.e., applying the above proof to the expanded-output channels), so $\vec{\mathcal{M}}$ is also $(\alpha,\sum_{j=1}^{J}\epsilon_{j},\omega)$-ms-Ave-SRPP.
Case 2: $\Xi\in\{\textup{Joint},\textup{ms-Joint}\}$. Fix $x$ in the support of $\mu_{s_{i}}^{\theta}$ and $\mu_{s_{j}}^{\theta}$. Recall that Joint-SRD can be represented as the Rényi divergence of a randomized sliced channel. Define the (one-shot) slicing operator
$\mathtt{Slice}(z):=(U,\langle z,U\rangle),\qquad U\sim\omega,$
where $U$ is independent of all other randomness. For a mechanism $\mathcal{M}$, write $\mathcal{N}:=\mathtt{Slice}\circ\mathcal{M}$. Then, by the definition/representation of Joint-SRD ( Definition 4.3 and Appendix B.3),
$\mathtt{JSD}^{\omega}_{\alpha}\bigl(\Pr(\mathcal{M}(X)\mid s_{i},\theta)\,\big\|\,\Pr(\mathcal{M}(X)\mid s_{j},\theta)\bigr)$ (78)
$=\mathtt{D}_{\alpha}\bigl(\Pr(\mathcal{N}(X)\mid s_{i},\theta)\,\big\|\,\Pr(\mathcal{N}(X)\mid s_{j},\theta)\bigr).$
@@
Now define, for each $j$, the sliced channel $\mathcal{N}_{j}:=\mathtt{Slice}\circ\mathcal{M}_{j}$. Since $\mathcal{M}_{j}$ is $(\alpha,\epsilon_{j},\omega)$-Joint-SRPP, (78) implies
$\mathtt{D}_{\alpha}\bigl(\Pr(\mathcal{N}_{j}(X)\mid s_{i},\theta)\,\big\|\,\Pr(\mathcal{N}_{j}(X)\mid s_{j},\theta)\bigr)\leq\epsilon_{j}.$
Consider the componentwise slicing of the composed mechanism: sample $U_{1},\dots,U_{J}\overset{\text{i.i.d.}}{\sim}\omega$ independently of everything, and define
@@
$\mathtt{Slice}^{J}\circ\vec{\mathcal{M}}(X):=\bigl((U_{1},\langle\mathcal{M}_{1}(X),U_{1}\rangle),\dots,(U_{J},\langle\mathcal{M}_{J}(X),U_{J}\rangle)\bigr).$
Equivalently, $\mathtt{Slice}^{J}\circ\vec{\mathcal{M}}(X)=(\mathcal{N}_{1}(X),\dots,\mathcal{N}_{J}(X))$. By the assumed independence of the primitive randomness across $j$ (and independence of $U_{j}$), the conditional laws under $s_{i}$ and $s_{j}$ factorize:
$\Pr\bigl((\mathcal{N}_{1}(x),\dots,\mathcal{N}_{J}(x))\mid s_{i},\theta\bigr)=\bigotimes_{j=1}^{J}\Pr\bigl(\mathcal{N}_{j}(x)\mid s_{i},\theta\bigr),$
and similarly under $s_{j}$. Therefore, by Rényi tensorization,
@@
$\mathtt{D}_{\alpha}\Bigl(\Pr\bigl(\mathtt{Slice}^{J}\circ\vec{\mathcal{M}}(x)\mid s_{i},\theta\bigr)\,\Big\|\,\Pr\bigl(\mathtt{Slice}^{J}\circ\vec{\mathcal{M}}(x)\mid s_{j},\theta\bigr)\Bigr)$
$=\sum_{j=1}^{J}\mathtt{D}_{\alpha}\Bigl(\Pr\bigl(\mathcal{N}_{j}(x)\mid s_{i},\theta\bigr)\,\Big\|\,\Pr\bigl(\mathcal{N}_{j}(x)\mid s_{j},\theta\bigr)\Bigr)\leq\sum_{j=1}^{J}\epsilon_{j}.$
Finally, by applying the representation (78) to the composed mechanism (with $\mathtt{Slice}^{J}$ as the slicing operator on $(\mathbb{R}^{d})^{J}$), the left-hand side is exactly the Joint-SRD of $\vec{\mathcal{M}}$ under the same slice profile $\omega$. Hence $\vec{\mathcal{M}}$ is $(\alpha,\sum_{j=1}^{J}\epsilon_{j},\omega)$-Joint-SRPP.
For ms-Joint-SRPP, the same argument applies after folding the subsampling randomness into each $\mathcal{M}_{j}$ (equivalently, applying the above proof to the corresponding expanded-output channels).
@@
Appendix W Experiment Details